# Explaining the Behavior of Black-Box Prediction Algorithms with Causal Learning

### v5

@@ 1 Introduction
To obtain impartial audits, it is desirable to have independent organizations or teams evaluate proposed prediction algorithms. This often means such audits must be performed without access to the inner workings of the target algorithm, due to intellectual property concerns. In such cases, post-hoc explainability methods that do not require access to the inner workings of the target algorithm are most appropriate. In this post-hoc setting, existing ML explainability approaches that involve retraining, modifying, or building newly explainable algorithms (Alvarez Melis & Jaakkola, 2018; Koh et al., 2020), or accessing intermediate layers (Kim et al., 2018; Bahadori & Heckerman, 2020; Yeh et al., 2020), are not directly applicable. “Feature importance” measures are popular; these aim to capture some generalized notion of association between each feature and a model’s prediction output (or prediction error), but do not generally differentiate between features that are merely associated with the model prediction versus those that have a causal effect on the output. We discuss this distinction and, more broadly, the importance of causal explainability in Section 2; see also Shin (2021) and Moraffah et al. (2020). Existing causality-aware feature importance approaches (Friedman, 2001; Loftus et al., 2023; Goldstein et al., 2015; Janzing et al., 2020) have important shortcomings in this post-hoc setting: (i) the explanatory features are micro-level inputs into the target prediction model, e.g., image pixels, rather than interpretable macro-level features that are more useful for understanding how to possibly change the algorithm’s behavior, and (ii) existing approaches assume there exists no unmeasured confounding between features and target model predictions, which fails to hold when the explanatory units are macro-level variables.
@@ 2 Causal Explanation
On Woodward’s counterfactual theory of causal explanation, explanations answer what-would-have-been-different questions. Specifically, the relevant counterfactuals describe the outcomes of interventions or manipulations. $X$ helps explain $Y$ if, under suitable background conditions, some intervention on $X$ produces a change in the distribution of $Y$. (Here we presume the object of explanation to be the random variable $Y$, not a specific value or event $Y=y$. That is, we choose to focus on type-level explanation rather than token-level explanations of particular events.) This perspective has strong connections to the literature on causal models in artificial intelligence and statistics (Spirtes et al., 2000; Pearl, 2009; Peters et al., 2017). A causal model for outcome $Y$ precisely stipulates how $Y$ would change under various interventions. So, to explain black-box algorithms we endeavor to build causal models for their behaviors. We propose that such causal explanations can be useful for algorithm evaluation and informing decision-making. In contrast, purely associational measures will be symmetric, include potentially irrelevant information, and fail to support (interventionist) counterfactual reasoning.
@@ 3 Causal Modeling
There are a variety of algorithms for causal structure learning, but what most approaches share is that they exploit patterns of statistical constraints implied by distinct causal models to distinguish among candidate graphs. One paradigm is constraint-based learning, which will be our focus. In constraint-based learning, the aim is to select a causal graph or set of causal graphs consistent with observed data by directly testing a sequence of conditional independence hypotheses. Distinct models will imply distinct patterns of conditional independence, and so by rejecting (or failing to reject) a collection of independence hypotheses, one may narrow down the set of models consistent with the data. (These methods typically rely on some version of the faithfulness assumption (Spirtes et al., 2000), which stipulates that all observed independence constraints correspond to missing edges in the graph.) For example, a classic constraint-based method is the PC algorithm Spirtes et al. (2000), which aims to learn an equivalence class of DAGs by starting from a fully-connected model and removing edges when conditional independence constraints are discovered via statistical tests. Since multiple DAGs may imply the same set of conditional independence constraints, PC estimates a CPDAG (completed partial DAG), a mixed graph with directed and undirected edges that represents a Markov equivalence class of DAGs. (Two graphs are called Markov equivalent if they imply the same conditional independence constraints.) Variations on the PC algorithm and related approaches to selecting CPDAGs have been thoroughly studied in the literature (Colombo & Maathuis, 2014; Cui et al., 2016).
[…]
@@
Figure 2: The PAG representation of the process above, with vertices $(Z_{1},Z_{2},Z_{4},Z_{5},\widehat{Y})$.
@@ 3.3 Limitations Of Existing Methods
Given this emphasis on macro-level explanatory variables and the attendant issue of unmeasured confounding, existing approaches to explainability have important limitations. One popular class of methods are “perturbation-based” approaches (Chattopadhyay et al., 2019; Schwab & Karlen, 2019; Selvaraju et al., 2017; Simonyan et al., 2013; Shrikumar et al., 2017), which perturb each element $X_{i}$, holding the remaining elements $X_{-i}$ fixed and observe the change in $\widehat{Y}$. Since the relationships among the $X$s and $Z$s are complex and unknown, these perturbation-based quantities cannot be straightforwardly translated into explanations as a function of macro-level explanatory units ($Z$s). Some causality-aware feature importance approaches, such as Partial Dependence Plots (Friedman, 2001), Causal Dependance Plots (Loftus et al., 2023), Individual Conditional Expectation plots (Goldstein et al., 2015), and modified causal variations of classical Shapley values (Janzing et al., 2020; Jung et al., 2022; Heskes et al., 2020) may be applied (or adapted) to provide explanatory information at the level of macro variables. However, these approaches assume there is no unmeasured confounding for the chosen set of explanatory features. Ultimately, these approaches will estimate quantities that adjust for (condition on) some subset of the observed features, but in the presence of unmeasured confounding there may be no observed subset that is valid. This can lead to substantially incorrect “explanatory” conclusions.
@@
Detailed calculations may be found in the Appendix. Now, if $z_{1}=\mathbb{E}[Z_{1}]$, then the incorrect Shapley value will be 00, while the true Shapley value will be $\frac{1}{2}\Big{\{}\mu_{0}\mathbb{E}[Z_{1}]\mathbb{E}[Z_{3}]-\mu_{0}\mathbb{E}[Z_{1}Z_{3}]\Big{\}}$, which can be non-zero. This illustrates a key shortcoming of Shapley values in this setting: if there is unmeasured confounding, even causally-informed Shapley values may lead to misleading conclusions about feature importance. There are promising recent proposals combining Shapley with identification theory for mixed graphs representing confounded variable sets (Jung et al., 2022), but in the setting considered here, the correct graph is not known.
@@ 4 Explaining Black-Box Predictions
with all errors $\epsilon_{i},\nu_{j}$ mutually independent. The marginal distribution over variables $(Z,\widehat{Y})$ satisfies the Markov condition wrt a MAG $\mathcal{M}$ (a projection of the DAG $\operatorname*{\mathcal{G}}$ onto this subset of vertices). The Markov equivalence class of $\mathcal{M}$ is represented by PAG $\mathcal{P}$ (Zhang, 2008a).
Our proposal is to estimate a causal PAG $\widehat{\mathcal{G}}$ over the variable set $V=(Z,\widehat{Y})$, with the minimal background knowledge imposed that the predicted outcome $\widehat{Y}$ is a causal non-ancestor of $Z$ (there is no directed path from $\widehat{Y}$ into any element of $Z$). This knowledge simply reflects the fact that the output of a prediction algorithm trained on an image does not cause the content of that image. Assuming faithfulness, the FCI algorithm is a consistent estimator of $\mathcal{P}$ (Spirtes et al., 2000; Zhang, 2008b).
@@
[…]
@@ 5 A Simulation Study
Figure 7: (a) A causal diagram representing the true data generating process. Here $Y$ denotes the true binary label and $C,T,H,$ and $V$ indicate the presence of circles, triangles, horizontal bars, and vertical bars in the images. $\widehat{Y}$ denotes the output of the combined neural network and logistic regression classifier. (b) The PAG learned using FCI over $(C,T,V,\widehat{Y})$.
@@
The model performs reasonably well: 77.89% accuracy on a set of 20,000 images. Our interpretable features $Z$ are indicators for the presence of the various shapes in the image. Since the true underlying behavior is causally determined by $V$ and $H$, we expect $V$ and $H$ to be “important” for the predictions $\widehat{Y}$, but due to the mutual dependence the other shapes are also highly correlated with $\widehat{Y}$. Moreover, we mimic a setting where $H$ is (wrongfully) omitted from the set of candidate interpretable features; in practice, the feature set proposed by domain experts or available in the annotated data will typically exclude various relevant determinants of the underlying outcome. In that case, $H$ is an unmeasured confounder. Applying the PAG-estimation algorithm FCI to variable set $(C,T,V,\widehat{Y})$ we learn the structure in Figure 7(b), which indicates correctly that $V$ is a (possible) cause, but that $T$ is not: $T$ is associated with the prediction outcomes only due to an unmeasured common cause. (FCI incorporates the user-specified background knowledge that $\widehat{Y}$ is not a cause of any image features.) [^1] Throughout, we use FCI as implemented in the command-line interface to the TETRAD freeware: https://github.com/cmu-phil/tetrad. This simple example illustrates how the estimated PAG using incomplete interpretable features can be useful. We disentangle mere statistical associations from (potentially) causal relationships, in this case indicating correctly that interventions on $V$ may make a difference to the prediction algorithm’s behavior but interventions on $T$ would not, and moreover that there is a potentially important cause of the output ($H$) that has been excluded from our set of candidate features, as evident from the bidirected edge $T\leftrightarrow\widehat{Y}$ learned by FCI. Existing approaches that do not properly account for unmeasured confounding cannot distinguish between such features and would incorrectly identify $T$ as “important” for explaining $\widehat{Y}$.
@@ 6.2 Structure Learning
[…]
@@ 6.3 Results
$\widehat{Y}$hroat ColorWing ShapeWing PatternUnderparts ColorUpper Tail Color
@@ 7.3 Selecting or Constructing Interpretable Features
In our experiments we use hand-crafted interpretable features that are available in the form of annotations with the data. Annotations are not always available in applications. In such settings, one approach would be to manually annotate the raw data with expert evaluations, e.g., when clinical experts annotate medical images with labels of important features (e.g. “tumor in the upper lobe,” “effusion in the lung,” etc).
Alternatively, one may endeavor to extract interpretable features automatically from the raw data. Unsupervised learning techniques may be used in some contexts to construct features, though in general there is no guarantee that these will be substantively (e.g., clinically) meaningful or correspond to manipulable elements of the domain. We explored applying some recent proposals for automatic feature extraction methods to our problem, but unfortunately, none of the available techniques proved appropriate, for reasons we discuss.
@@
Alternative approaches to automatic feature selection also proved problematic for our purposes. In Lopez-Paz et al. (2017), the authors train an object detection network that serves as a feature extractor for future images, and then they use these extracted features to infer directions of causal effects. However, this approach relies on the availability of training data with a priori associated object categories: the Pascal Visual Object Classes, which are labels for types of objects (e.g., aeroplane, bicycle, bird, boat, etc.) (Everingham et al., 2015). This approach is thus not truly data-driven and relies on auxiliary information that is not generally available. In another line of work, various unsupervised learning methods (based on autoencoders or variants of independent component analysis) are applied to extract features from medical images to improve classification performance (Arevalo et al., 2015; Sari & Gunduz-Demir, 2018; Jiang et al., 2018). These approaches are more data-driven but do not produce interpretable features – they produce a compact lower-dimensional representation of the pixels in an abstract vector space, which does not necessarily map onto something a user may understand or recognize.

### v6

@@ 1 Introduction
To obtain impartial audits, it is desirable to have independent organizations or teams evaluate proposed prediction algorithms. This often means such audits must be performed without access to the inner workings of the target algorithm, due to intellectual property concerns. In such cases, post-hoc explainability methods that do not require access to the inner workings of the target algorithm are most appropriate. In this post-hoc setting, existing ML explainability approaches that involve retraining, modifying, or building newly explainable algorithms (Alvarez Melis & Jaakkola, 2018; Koh et al., 2020), or accessing intermediate layers (Kim et al., 2018; Bahadori & Heckerman, 2021; Yeh et al., 2020), are not directly applicable. “Feature importance” measures are popular; these aim to capture some generalized notion of association between each feature and a model’s prediction output (or prediction error), but do not generally differentiate between features that are merely associated with the model prediction versus those that have a causal effect on the output. Distinguishing whether some variable is a cause of model outputs rather than only associated due to confounding is crucial to understanding whether manipulations of that feature will make any difference to the prediction model output; intervening on variables associated only due to confounding will have no consequence and provides no explanation of the outcome. We discuss this distinction and, more broadly, the importance of causal explainability in Section 2; see also Shin (2021) and Moraffah et al. (2020). Existing causality-aware feature importance approaches (Friedman, 2001; Loftus et al., 2024; Goldstein et al., 2015; Janzing et al., 2020) have important shortcomings in this post-hoc setting: (i) the explanatory features are micro-level inputs into the target prediction model, e.g., image pixels, rather than interpretable macro-level features that are more useful for understanding how to possibly change the algorithm’s behavior, and (ii) existing approaches assume there exists no unmeasured confounding between features and target model predictions, which fails to hold when the explanatory units are macro-level variables. Throughout, we use “micro-level” to refer to the raw features directly input into the prediction model, which are typically pixel or voxel values in image analysis and may be words or tokens in text analysis. “Macro-level” features then correspond to (observed or latent) abstractions of the raw features that are independently manipulable, also described as causal representations or disentangled representations (Schölkopf et al., 2021).
@@ 2 Causal Explanation
On Woodward’s counterfactual theory of causal explanation, explanations answer what-would-have-been-different questions. Specifically, the relevant counterfactuals describe the outcomes of interventions or manipulations. $X$ helps explain $Y$ if, under suitable background conditions, some intervention on $X$ produces a change in the distribution of $Y$. (Here we presume the object of explanation to be the random variable $Y$, not a specific value or event $Y=y$. That is, we choose to focus on type-level explanation rather than token-level explanations of particular events. See Beckers (2022) for related discussion focused on causal explanation at the event level.) This perspective has strong connections to the literature on causal models in artificial intelligence and statistics (Spirtes et al., 2000; Pearl, 2009; Peters et al., 2017). A causal model for outcome $Y$ precisely stipulates how $Y$ would change under various interventions. So, to explain black-box algorithms we endeavor to build causal models for their behaviors. We propose that such causal explanations can be useful for algorithm evaluation and informing decision-making. In contrast, purely associational measures will be symmetric, include potentially irrelevant information, and fail to support (interventionist) counterfactual reasoning.
@@ 3 Causal Modeling
There are a variety of algorithms for causal structure learning, but what most approaches share is that they exploit patterns of statistical constraints implied by distinct causal models to distinguish among candidate graphs. One paradigm is constraint-based learning, which will be our focus. In constraint-based learning, the aim is to select a causal graph or set of causal graphs consistent with observed data by directly testing a sequence of conditional independence hypotheses. Distinct models will imply distinct patterns of conditional independence, and so by rejecting (or failing to reject) a collection of independence hypotheses, one may narrow down the set of models consistent with the data. (These methods typically rely on some version of the faithfulness assumption (Spirtes et al., 2000), which stipulates that all observed independence constraints correspond to missing edges in the graph.) For example, a classic constraint-based method is the PC algorithm (Spirtes et al., 2000), which aims to learn an equivalence class of DAGs by starting from a fully-connected model and removing edges when conditional independence constraints are discovered via statistical tests. Since multiple DAGs may imply the same set of conditional independence constraints, PC estimates a CPDAG (completed partial DAG), a mixed graph with directed and undirected edges that represents a Markov equivalence class of DAGs. (Two graphs are called Markov equivalent if they imply the same conditional independence constraints.) Variations on the PC algorithm and related approaches to selecting CPDAGs have been thoroughly studied in the literature (Colombo & Maathuis, 2014; Cui et al., 2016).
In settings with unmeasured (latent) confounding variables, it is typical to study graphs with bidirected edges to represent dependence due to confounding. For example, a partial ancestral graph (PAG) (Zhang, 2008a) is a graphical representation which includes directed edges ($X\to Y$ means $X$ is a causal ancestor of $Y$), bidirected edges ($X\leftrightarrow Y$ means $X$ and $Y$ are both caused by some unmeasured common factor(s), e.g., $X\leftarrow U\rightarrow Y$), and partially directed edges ($X\operatorname{\hbox{$\circ$}\kern-1.5pt\hbox{$\rightarrow$}}Y$ or $X\operatorname{\hbox{$\circ$}\kern-1.2pt\hbox{$--$}\kern-1.5pt\hbox{$\circ$}}Y$) where the circle marks indicate ambiguity about whether the endpoints are arrowheads or tails. Generally, PAGs may also include additional edge types to represent selection bias, but this is irrelevant for our purposes here. PAGs inherit their causal interpretation by encoding the commonalities among a set of underlying causal DAGs with latent variables. A bit more formally, a PAG represents an equivalence class of maximal ancestral graphs (MAGs), which encode the independence relations among observed variables when some variables are unobserved (Richardson & Spirtes, 2002; Zhang, 2008a). The FCI algorithm (Spirtes et al., 2000; Zhang, 2008b) is a well-known constraint-based method, which uses sequential tests of conditional independence to select a PAG from data. Similarly to the PC algorithm, FCI begins by iteratively deleting edges from a fully-connected graph by a sequence of conditional independence tests. However, the space of possible models (mixed graphs with multiple types of edges) is much greater for FCI as compared with PC, so the rules linking patterns of association to possible causal structures are more complex. Though only independence relations among observed variables are testable, it can be shown formally that certain patterns of conditional independence among observed variables are incompatible with certain latent structures (assuming faithfulness), so some observed patterns will rule out unmeasured confounding and other observed patterns will on the contrary suggest the existence of unmeasured confounders. Details of the FCI algorithm may be found in the literature (Zhang, 2008b; Colombo et al., 2012). Variations on the FCI algorithm and alternative PAG-learning procedures have also been studied (Colombo et al., 2012; Claassen & Heskes, 2012; Ogarrio et al., 2016; Tsirlis et al., 2018).
@@ 3.3 Limitations Of Existing Methods
Given this emphasis on macro-level explanatory variables and the attendant issue of unmeasured confounding, existing approaches to explainability have important limitations. One popular class of methods are “perturbation-based” approaches (Chattopadhyay et al., 2019; Schwab & Karlen, 2019; Selvaraju et al., 2017; Simonyan et al., 2013; Shrikumar et al., 2017; Chen et al., 2024), which perturb each element $X_{i}$, holding the remaining elements $X_{-i}$ fixed and observe the change in $\widehat{Y}$. Since the relationships among the $X$s and $Z$s are complex and unknown, these perturbation-based quantities cannot be straightforwardly translated into explanations as a function of macro-level explanatory units ($Z$s). Some causality-aware feature importance approaches, such as Partial Dependence Plots (Friedman, 2001), Causal Dependance Plots (Loftus et al., 2024), Individual Conditional Expectation plots (Goldstein et al., 2015), and modified causal variations of classical Shapley values (Janzing et al., 2020; Jung et al., 2022; Heskes et al., 2020) may be applied (or adapted) to provide explanatory information at the level of macro variables. However, these approaches assume there is no unmeasured confounding for the chosen set of explanatory features. Ultimately, these approaches will estimate quantities that adjust for (condition on) some subset of the observed features, but in the presence of unmeasured confounding there may be no observed subset that is valid. This can lead to substantially incorrect “explanatory” conclusions.
@@
Detailed calculations may be found in the Appendix. Now, if $z_{1}=\mathbb{E}[Z_{1}]$, then the incorrect Shapley value will be 0, while the true Shapley value will be $\frac{1}{2}\Big{\{}\mu_{0}\mathbb{E}[Z_{1}]\mathbb{E}[Z_{3}]-\mu_{0}\mathbb{E}[Z_{1}Z_{3}]\Big{\}}$, which can be non-zero. This illustrates a key shortcoming of Shapley values in this setting: if there is unmeasured confounding, even causally-informed Shapley values may lead to misleading conclusions about feature importance. There are promising recent proposals combining Shapley with identification theory for mixed graphs representing confounded variable sets (Jung et al., 2022), but in the setting considered here, the correct graph is not known.
@@ 4 Explaining Black-Box Predictions
[…]
We primarily use the FCI algorithm in our experiments, but alternative PAG-learning discovery algorithms may be also be used. FCI selects a PAG by executing the following steps: (1) beginning with a complete graph over the vertices, a series of conditional independence hypotheses are tested for every pair of verticies and every possible conditioning set. When an independence is discovered between two vertices, the edge between them is removed and the relevant conditioning set is recorded (the separating set). (2) The procedure then orients some triples of variables as colliders using the so-called “collider rule”: a triple $Z_{i}\operatorname{\hbox{$\circ$}\kern-1.2pt\hbox{$--$}\kern-1.5pt\hbox{$\circ$}}Z_{j}\operatorname{\hbox{$\circ$}\kern-1.2pt\hbox{$--$}\kern-1.5pt\hbox{$\circ$}}Z_{k}$ with $Z_{i}$ and $Z_{k}$ not adjacent is oriented $Z_{i}\operatorname{\hbox{$\circ$}\kern-1.5pt\hbox{$\rightarrow$}}Z_{j}\operatorname{\hbox{$\leftarrow$}\kern-1.5pt\hbox{$\circ$}}Z_{k}$ only if $Z_{j}$ is not in the separating set for $Z_{i},Z_{k}$. The result at this stage is a partially-oriented graph. (3) Some additional independence tests are carried out based on the graph structure (see detailed discussion in Colombo et al. (2012)) and the collider rule is repeated. (4) Finally, some additional orientations may follow from the orientations already determined, using provably complete graphical rules that are applied exhuastively to the partially oriented graph (Zhang, 2008b). Assuming faithfulness, the FCI algorithm is an asymptotically correct estimator of $\mathcal{P}$ (Spirtes et al., 2000; Zhang, 2008b; Colombo et al., 2012). Simulations previously reported indicate that when the ground truth is sparse and the sample size is sufficiently large, FCI performs well and recovers a structure close to the true PAG. However, every conditional independence test performed has some chance of error, and compounding errors may lead to the incorrect structure especially when sample size is small or the truth is dense. Though it is very difficult to exactly recover the correct graph with any finite sample size, simulation studies show that graphs with good precision can in many cases be estimated from on the order of a few thousand samples (Colombo et al., 2012; Ogarrio et al., 2016).
Additional background knowledge may also be imposed if it is known, for example, that none of the elements of $Z$ may cause each other (they are only dependent due to latent common causes) or there are groups of variables which precede others in a causal ordering. If in the estimated graph $\widehat{\mathcal{G}}$, there is a directed edge $Z_{i}\rightarrow\widehat{Y}$, then $Z_{i}$ is a cause (definite causal ancestor) of the prediction, if instead there is a bidirected edge $Z_{i}\leftrightarrow\widehat{Y}$ then $Z_{i}$ is not a cause of the prediction but they are dependent due to common latent factors, and if $Z_{i}\operatorname{\hbox{$\circ$}\kern-1.5pt\hbox{$\rightarrow$}}\widehat{Y}$ then $Z_{i}$ is a possible cause (possible ancestor) of the prediction but unmeasured confounding cannot be ruled out. These edge types are summarized in Table 1. The reason it is important to search for a PAG and not a DAG (or equivalence class of DAGs) is that $Z$ will in general not include all possibly relevant variables.
@@
$Z_{i}\operatorname{\hbox{$\circ$}\kern-1.5pt\hbox{$\rightarrow$}}\widehat{Y}$ Either $Z_{i}$ is a cause of $\widehat{Y}$ or there is unmeasured confounding, or both
@@ 5 A Simulation Study
[…]
@@
[…]
@@ 6.2 Structure Learning
Structure learning methods such as FCI produce a single estimated PAG as output. In applications, it is common to repeat the graph estimation on multiple bootstraps or subsamples of the original data, in order to control false discovery rates or to mitigate the cumulative effect of statistical errors in the independence tests (Stekhoven et al., 2012). We create 50 bootstrapped replicates of the bird image dataset and 100 replicates of the X-ray dataset. We run FCI on each replicate with independence test rejection threshold (a tuning parameter) set to $\alpha=.05$ and $\alpha=.01$ for the birds and X-ray datasets, respectively, with the knowledge constraint imposed that outcome $\widehat{Y}$ cannot cause any of the interpretable features. Here FCI is used with the $\chi^{2}$ independence test, and we limit the maximum conditioning set size to 4 for computational tractability in the birds dataset. (No limit is necessary in the smaller X-ray dataset.) As a sensitivity analysis, we repeat both experiments using a different PAG-learning procedure, a hybrid algorithm that combines the greedy score-based procedure GRaSP (Lam et al., 2022) with FCI. The details are described in Appendix C.
@@ 6.3 Results
Our sensitivity analysis using an alternative PAG-learning procedure, GRaSP-FCI, produces similar results. In the X-ray experiment, the final ordering of features by edge stability GRaSP-FCI is the same as the ordering produced by FCI. In the bird classification experiment, there are some substantive differences: a few features ranked highly by FCI are ranked low by GRaSP-FCI and vice versa, though for most features the conclusions are concordant. Crown Color and Size are ranked highest. The full results are presented in Appendix C.
@@
$\widehat{Y}$Throat ColorWing ShapeWing PatternUnderparts ColorUpper Tail Color
@@ 7.3 Selecting or Constructing Interpretable Features
In our experiments we use hand-crafted interpretable features that are available in the form of annotations with the data. Annotations are not always available in applications. In such settings, one approach would be to manually annotate the raw data with expert evaluations, e.g., when clinical experts annotate medical images with labels of important features (e.g. “tumor in the upper lobe,” “effusion in the lung,” etc). In settings where annotations are available only on a small subset of examples, semi-supervised learning techniques developed in recent years may be useful for generating “pseudo-labels,” and this may make the approach described here feasible even in settings with only a few annotated examples. Pipelines such as FixMatch (Sohn et al., 2020) use the predictions from a supervised model trained on annotated data combined with multiple “augmented” versions of unlabeled images that are expected to having matching labels to generate pseudo-labels at scale with excellent predictive performance.
Alternatively, one may endeavor to extract interpretable features automatically from the raw unlabeled data. Unsupervised learning techniques may be used in some contexts to construct features, though in general there is no guarantee that these will be substantively (e.g., clinically) meaningful or correspond to manipulable elements of the domain. We explored applying some recent proposals for automatic feature extraction methods to our problem, but unfortunately, none of the available techniques proved appropriate, for reasons we discuss.
@@
Alternative approaches to automatic feature selection also proved problematic for our purposes. In Lopez-Paz et al. (2017), the authors train an object detection network that serves as a feature extractor for future images, and then they use these extracted features to infer directions of causal effects. However, this approach relies on the availability of training data with a priori associated object categories: the Pascal Visual Object Classes, which are labels for types of objects (e.g., aeroplane, bicycle, bird, boat, etc.) (Everingham et al., 2015). This approach is thus not truly data-driven and relies on auxiliary information that is not generally available. In another line of work, various unsupervised learning methods (based on autoencoders or variants of independent component analysis) are applied to extract features from medical images to improve classification performance (Arevalo et al., 2015; Sari & Gunduz-Demir, 2018; Jiang et al., 2018). These approaches are more data-driven but do not produce interpretable features – they produce a compact lower-dimensional representation of the pixels in an abstract vector space, which does not necessarily map onto something a user may understand or recognize. Recently, there has been interest in using large language models (LLMs) as annotators for unlabeled data, either alone or in collaboration with humans (Wang et al., 2024; Tang et al., 2024; Beck et al., 2025). This technology and its use for annotation is still in its development, but there is growing evidence that with careful engineering and human input, LLM-augmented annotations may cut costs for generating annotations at scale but also face limitations that are important to better understand.
@@
Finally, we note that macro-level feature annotations, whether they are derived from human judgment or statistical models, may contain errors. When labels are “noisy” or are corrupted by measurement error, this can degrade the performance of causal discovery algorithms. Some approaches have been proposed for mitigating the effect of measurement error on constraint-based algorithms (Blom et al., 2018), but in general this is a challenging and mostly open problem.